a. Two hunters average one hit per three shots. They both see a fox and fire at the same moment. What is the probability that at least one of the hunters will hit the fox?

b. Solve the same problem with n numbers, each of whom has the same probability as the hunters above.
Answer #1

a)

here probability that a hunter hits the fox =1/3

and probability that a hunter does not hits the fox=1-1/3=2/3

hence P( at least one of the hunters will hit the fox )=1-P(none of hunter hits the fox)

=1-P(first does not hit and second does not hit)=1-(2/3)*(2/3)=5/9

b)

P( at least one of the hunters will hit the fox )=1-P(none of hunter hits the fox)

probability =1-(2/3)n